how many grams of oxygen are produced when 0.250 mol NO2 is formed?
1
Expert's answer
2017-02-15T07:23:43-0500
Solution:
M (O2) = 32 g/ mol
n (NO2) = 0.250 mol
Looking at the reaction you see that there is x times as much mol NO2.
Need a reaction in which NO2 is formed (A(s)--->xNO2(g)+ yC(g))
O2= 0.25/x = 0.25/x mol.
The answer in grams:
0.25/x mol * 32g / mol= (8 / x) grams
Answer: (8 / x) grams